OpenStudy (anonymous):

show checking step& other work please 6x - 11 = 16 + 9x

OpenStudy (bahrom7893):

6x - 11 = 16 + 9x 6x - 9x = 16 + 11 -3x = 27 x = -9

OpenStudy (anonymous):

9x -6x = -11 -16 3x = -27 =>x = -27/3 = -9